# Two-dimensional fractional Brownian motion: Analysis in time and frequency domains

### v1

@@ I Introduction
The study of anomalous diffusion phenomena has emerged as a critical area of research in diverse scientific fields, from cellular biophysics to financial mathematics [34, 24, 31, 47, 43, 37]. While standard Brownian motion adequately describes normal diffusion processes characterized by a mean-square displacement that increases linearly with time [14], multiple complex systems exhibit a non-linear scaling that demands a more sophisticated mathematical framework [34]. Fractional Brownian motion (fBm), formalized by Mandelbrot and van Ness extends the classical Brownian motion model by introducing long-range dependence through the Hurst parameter $H$ [30]. However, when considering multidimensional systems, the scalar parameter $H$ can be insufficient to capture the complex cross-dependencies and anisotropic scaling behaviors observed in empirical data. Multidimensional fBm or vector fBm [33, 20, 28, 1], thus, represents an advancement in the modeling of such anomalous diffusion processes in multidimensional settings.
@@
In this article, we propose the construction of a two-dimensional fractional Brownian motion (2D fBm) by introducing a dependence on the underlying noise in the time representation of the process. This construction explicitly considers cross-dependencies and anisotropic scaling by using a matrix-valued Hurst operator that allows direction-dependent scaling properties. We examine the theoretical properties of the proposed construction and derive the auto- and cross-covariance structure of the process and its increments. In addition, we calculate the power spectral density (PSD) of these processes [40], a fundamental measure often used in the characterization of time series across different disciplines [18, 23, 13, 15, 5]. At last, we present numerical simulations that showcase and validate our analytical results. Although multidimensional fBm with dependent components has been investigated in the existing literature[1, 32], our model has the advantage of having a relatively simple construction, based on the correlated Brownian motions used in its integral representation. The new approach uniquely allows for any range of model parameters, a significant advantage over the previous constructions[1]. Consequently, this simplified construction enables a thorough examination of the process characteristics, particularly with respect to its auto- and cross-covariance structures in both the time and frequency domains.
@@ II Model
In both cases, $\mathrm{d}\tilde{W}_{1}(t),\mathrm{d}\tilde{W}_{2}(t)$ are $\rho$-correlated Gaussian noises on the real line, $H_{1},H_{2}\in(0,1)$, and $\sigma_{1},\sigma_{2}>0$. The constants $a_{H_{1}}$, $a_{H_{2}}$, $a^{*}_{H_{1}}$, and $a^{*}_{H_{2}}$ are chosen in such a way that the processes $X_{j}(t)$ and $X^{*}_{j}(t)$ for $j=1,2$ have variances $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$ for $t=1$, respectively, that is, $\langle X_{j}^{2}(1)\rangle=\langle X^{*2}_{j}(1)\rangle=\sigma_{j}^{2},\textrm{ for }j=1,2$. Parameter $\rho$ will be referred to as correlation coefficient of the underlying noise. Let us note that for both processes $\mathbf{X}(t)$ and $\mathbf{X}^{*}(t)$, the marginals $X_{j}(t)$ and $X_{j}^{*}(t)$ ($j=1,2$) are fractional Brownian motions with corresponding Hurst parameters $H_{j}$ (cf. [30] for $\mathbf{X}(t)$ and [52] for $\mathbf{X}^{*}(t)$). The difference between the seemingly similar processes $\mathbf{X}(t)$ and $\mathbf{X}^{*}(t)$ lies in the cross-dependence structure between their components, specifically between $X_{1}(t)$ and $X_{2}(t)$, and between $X_{1}^{*}(t)$ and $X_{2}^{*}(t)$, respectively, which we discuss later in this section.
@@
$\mathbf{X}(t)=\int_{-\infty}^{\infty}\begin{bmatrix}\sigma_{1}a_{H_{1}}f_{+}(t;s,H_{1}-1/2)&0\\ 0&\sigma_{2}a_{H_{2}}f_{+}(t;s,H_{2}-1/2)\end{bmatrix}\begin{bmatrix}1&0\\ \rho&\sqrt{1-\rho^{2}}\end{bmatrix}\begin{bmatrix}\mathrm{d}W_{1}(s)\\ \mathrm{d}W_{2}(s)\end{bmatrix}$, (6)
@@
The process $\mathbf{X}(t)$ is a zero-mean Gaussian process and has a strong connection to the so-called operator fractional Brownian motion[42]. Such relation allows for analysing the covariance structure of the process $\mathbf{X}(t)$ (see Theorem 1 in Section III).
@@
Contrary to the causal construction presented in Definition 1, the well-balanced case adds to the previous case an anti-causal filtering of a Brownian motion. We now introduce such a process, together with the calculation of the normalization constant (Theorem 6 in Appendix B). The covariance structure of well-balanced 2D fBm is provided in Section III.
Defintition 2 (Well-balanced 2D fBm).
@@
for $s\in\mathbb{R},t\geq 0,j=1,2$, $\sigma_{1},\sigma_{2}>0$, and the constants $a^{*}_{H_{j}}$’s are given by
@@ III.1 Covariance structure of the process
The proof of this theorem is presented in Appendix A and Appendix C. ∎
@@ IV Spectral Content
The PSD matrix plays a central role in the analysis of multidimensional stochastic processes. This quantity generalizes the concept of a PSD to vector-valued processes. It is particularly valuable in multiple fields, including econometrics [17], structural engineering [11, 21], neuroscience [57], and signal processing [2], where complex systems exhibit interactions across multiple channels or variables. Since PSD is more frequently employed in the context of stationary processes, we begin this section by analyzing the increments of 2D fBm.
@@ IV.1 Spectral content of the increments
which is known as Wiener-Khinchin theorem[27]. Let us note, that for a $d$-dimensional process, the function $S_{\Delta\mathbf{Z}}$ is a $d\times d$ matrix with the diagonal elements corresponding to marginals and off-diagonal elements describing the, so called, cross power spectral density . To streamline the notation, we will use now $S^{\Delta}$ whenever PSD relates to the increments process, instead of using $S_{\Delta\mathbf{Z}}$.
@@
for $j,k=1,2$. Constants $a_{H_{j}}$ and $a^{*}_{H_{j}}$ are given in Eqs. (7) and (11), respectively. Then, for the increments of 2D fBm the power spectral density matrix $S^{\Delta}(f)$ is given by
@@
where $(x)_{+}\equiv\max\{x,0\},(x)_{-}\equiv\max\{-x,0\},D=H-0.5I_{2}$, and $\overline{C}$ denotes the element-wise complex conjugate of matrix $C$. Matrix $\tilde{C}$ is equal to $C_{\textrm{c}}$ or $C_{\textrm{wb}}$ depending if we consider causal or well-balanced 2D fBm. The components $\langle S^{\Delta}_{jk}\rangle$ can be thus expressed as
@@
The proof of this theorem is presented in Appendix D. ∎
@@
The direct derivation of matrix $C$ given in Theorem 3 be found in the proof of Theorem 1 given in Appendix A and Appendix C.
@@ IV.2 Spectral content of the process
The proof of this theorem is included in Appendix E. ∎
@@
$\left.-a_{jk}\left[\frac{\cot\left(\frac{\pi}{2}(H_{j}+H_{k})\right)}{\tilde{\omega}^{H_{j}+H_{k}+1}}+\frac{2+H_{j}+H_{k}-\cos\tilde{\omega}\sin\tilde{\omega}\cot\left(\frac{\pi}{2}(H_{j}+H_{k})\right)}{\tilde{\omega}^{H_{j}+H_{k}+2}}\right]\right\},$ (48)
@@
The proof of this theorem is based on the asymptotic expansion of the incomplete gamma function and is presented in details in Appendix F. ∎
@@
$\Re\langle S_{jk}(f,T)\rangle\sim\frac{\sigma_{j}\sigma_{k}\rho_{jk}}{f^{2}}T^{H_{j}+H_{k}-1}+\frac{2c_{H}\sigma_{j}\sigma_{k}}{f^{H_{j}+H_{k}+1}}+o(1).$ (49)
@@
$\Re\langle S_{jk}(f,T)\rangle\sim\frac{\sigma_{j}\sigma_{k}\rho_{jk}}{f^{H_{j}+H_{k}+1}}+\frac{2\sigma_{j}\sigma_{k}}{T^{1-H_{j}-H_{k}+1}}+o(T^{H_{j}+H_{k}-1}).$ (50)
@@ V.2 Autocovariance
[…]
Figure 4: Cross-covariance function depending on $H_{1},H_{2}$ and $\rho_{12}$ for $H_{1}=H_{2}$. (a) $H_{1}=H_{2}=0.2$, (b) $H_{1}=H_{2}=0.5$, and (c) $H_{1}=H_{2}=0.7$. The solid lines represent the causal version of the model and dashed lines the well-balanced case. Markers correspond to the estimated values of the cross-covariance function; circles correspond to the causal case, squares to the well-balanced case. Here, casual and well-balanced cases overlap, as for $H_{1}=H_{2}$ the model is time-reversible regardless of the definition (cf. Eq. (15)).
[…]
@@ V.3 Spectral content
[…]
Figure 6: Ensemble-averaged (cross-)PSD. Solid lines correspond to PSD estimated from the simulated trajectories, dashed lines to their asymptotics. (a) $\Re\langle S_{\mathbf{X},12}(f,T)\rangle$, $H_{1}=H_{2}$; (b) $\Re\langle S_{\mathbf{X},12}(f,T)\rangle$, $H_{1}\neq H_{2}$; and (c) $\Im\langle S_{\mathbf{X},12}(f,T)\rangle$, $H_{1}\neq H_{2}$. Since the real part for both causal and well-balanced case is the same, we plot only only one corresponding line to each set of parameteres. Conversely, the imaginary part for the well-balanced case is 0, so we plot only $\Im\langle S_{\mathbf{X},12}\rangle$ for the causal 2D fBm. In all the panels, $T=2^{16}=65,536$ and the lines are based on $N=5,000$ trajectories.
@@
In all panels, we see that the asymptotic behavior of the ensemble-averaged (cross-)PSD for increments behaves in a similar way in all cases, regardless if $H_{1}+H_{2}\gtrless 1$. The asymptotic behavior, as $f\to 0$ is proportional to $f^{1-H_{1}-H_{2}}$. Here, we see a quantitative difference of diverging or converging (cross-)PSD around 0 – if $H_{1}+H_{2}>1$ then the (cross-)PSD diverges (in 1D, it corresponds to one of the definitions of long memory [42]), while for $H_{1}+H_{2}$ the (cross-)PSD converges to 0.
[…]
Figure 7: Ensemble-averaged (cross-)PSD for increments. Continuous lines correspond to PSD estimated from the simulated trajectories’ increments, dashed lines to their asymptotics. (a) $\Re\langle S^{\Delta}_{12}(f,T)\rangle$, $H_{1}=H_{2}$; (b) $\Re\langle S^{\Delta}_{12}(f)\rangle$, $H_{1}\neq H_{2}$; and (c) $\Im\langle S^{\Delta}_{12}(f)\rangle$, $H_{1}\neq H_{2}$. Since the real part for both causal and well-balanced case is the same, we plot only only one corresponding line to each set of parameters. Conversely, the imaginary part for the well-balanced case is 0, so we plot only $\Im\langle S^{\Delta}_{12}(f)\rangle$ for the causal 2D fBm’s increments. In all the panels, $T=2^{16}=65,536$ and the lines are based on $N=5,000$ trajectories.

### v2

@@ Abstract
Two-dimensional fractional Brownian motion extends the idea of random motion by allowing both long-term memory and interdependence between different directions of movement. Unlike standard models, where the horizontal and vertical components are assumed to fluctuate independently, this process accounts for correlations and directional differences in scaling. We introduce these features through correlated underlying noises and the matrix-valued Hurst operator, which both make it possible for one direction to show, for example, slower subdiffusive dynamics while the other exhibits faster, superdiffusive behavior. We also show that, depending on how we choose to construct the process, the cross-correlations between components may be symmetric or asymmetric, leading to different temporal and spectral characteristics. The considered model is especially relevant for systems where anisotropy and interdependence cannot be ignored, such as the motion of particles in complex biological environments, the joint evolution of financial indices, or, simply, two correlated fractional Brownian motions.
@@ I Introduction
The study of anomalous diffusion phenomena has emerged as a critical area of research in diverse scientific fields, from cellular biophysics to financial mathematics [1, 2, 3, 4, 5, 6, 7]. While standard Brownian motion adequately describes normal diffusion processes characterized by a mean-square displacement that increases linearly with time [8], multiple complex systems exhibit a non-linear scaling that demands a more sophisticated mathematical framework [1]. Fractional Brownian motion (fBm) originates from the foundational work of Kolmogorov in the 1940s, who first formulated a Gaussian process with stationary increments and a power-law covariance structure[9]. It was later formalized by Mandelbrot and van Ness, who introduced the integral representation and the Hurst parameter $H$ as a descriptor of memory and anomalous scaling[10]. However, when considering multidimensional systems, the scalar parameter $H$ can be insufficient to capture the complex cross-dependencies and anisotropic scaling behaviors observed in empirical data. Multidimensional fBm or vector fBm [11, 12, 13, 14], thus, represents an advancement in the modeling of such anomalous diffusion processes in multidimensional settings.
@@
In this article, we propose the construction of a two-dimensional fractional Brownian motion (2D fBm) by introducing a dependence on the underlying noise in the time representation of the process. This construction explicitly considers cross-dependencies and anisotropic scaling by using a matrix-valued Hurst operator that allows direction-dependent scaling properties. Depending on the construction of the 2D fBm, we consider two versions of this process: causal and well-balanced. The well-balanced process is time-reversible, a property absent in the causal 2D fBm. As a result, these two formulations exhibit distinct correlation structures: symmetric cross-correlations for the well-balanced case and asymmetric cross-correlations for the causal case. We examine the theoretical properties of the two versions of 2D fBm and derive the auto- and cross-covariance structures of such process and their increments. In addition, we calculate the power spectral density (PSD) of these processes [44], a fundamental measure often used in the characterization of time series across different disciplines [45, 46, 47, 48, 49]. At last, we present numerical simulations that showcase and validate our analytical results. Although multidimensional fBm with dependent components has been investigated in the existing literature[14, 37], our proposition has the advantage of having a relatively simple construction, based on the correlated Brownian motions used in its integral representation. The new approach uniquely allows for any range of model parameters, a significant advantage over the previous constructions[14]. Consequently, this simplified construction enables a thorough examination of the process characteristics, particularly with respect to its auto- and cross-covariance structures in both the time and frequency domains.
@@ II Model
In both cases, $\mathrm{d}\tilde{W}_{1}(t),\mathrm{d}\tilde{W}_{2}(t)$ are $\rho$-correlated Gaussian noises on the real line, $H_{1},H_{2}\in(0,1)$, and $\sigma_{1},\sigma_{2}>0$. The constants $a_{H_{1}}$, $a_{H_{2}}$, $a^{*}_{H_{1}}$, and $a^{*}_{H_{2}}$ are chosen in such a way that the processes $X_{j}(t)$ and $X^{*}_{j}(t)$ for $j=1,2$ have variances $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$ for $t=1$, respectively, that is, $\langle X_{j}^{2}(1)\rangle=\langle X^{*2}_{j}(1)\rangle=\sigma_{j}^{2},\textrm{ for }j=1,2$. Parameter $\rho$ will be referred to as correlation coefficient of the underlying noise. Let us note that for both processes $\mathbf{X}(t)$ and $\mathbf{X}^{*}(t)$, the marginals $X_{j}(t)$ and $X_{j}^{*}(t)$ ($j=1,2$) are fractional Brownian motions with corresponding Hurst parameters $H_{j}$ (cf. [10] for $\mathbf{X}(t)$ and [50] for $\mathbf{X}^{*}(t)$). The difference between the seemingly similar processes $\mathbf{X}(t)$ and $\mathbf{X}^{*}(t)$ lies in the cross-dependence structure between their components, specifically between $X_{1}(t)$ and $X_{2}(t)$, and between $X_{1}^{*}(t)$ and $X_{2}^{*}(t)$, respectively, which we discuss in the next section.
@@
$\mathbf{X}(t)=\int_{-\infty}^{\infty}\begin{bmatrix}\sigma_{1}a_{H_{1}}f_{+}(s;t,H_{1}-1/2)&0\\ 0&\sigma_{2}a_{H_{2}}f_{+}(s;t,H_{2}-1/2)\end{bmatrix}\begin{bmatrix}1&0\\ \rho&\sqrt{1-\rho^{2}}\end{bmatrix}\begin{bmatrix}\mathrm{d}W_{1}(s)\\ \mathrm{d}W_{2}(s)\end{bmatrix}$, (6)
@@
The process $\mathbf{X}(t)$ is a zero-mean Gaussian process and has a strong connection to the so-called operator fractional Brownian motion[52, 53, 54, 55]. Such a relation allows for analyzing the covariance structure of the process $\mathbf{X}(t)$ (see Theorem 1 in section III).
@@
Contrary to the causal construction presented in Definition 1, the well-balanced case adds to the previous case an anti-causal filtering of a Brownian motion. We now introduce such a process, together with the calculation of the normalization constant (Theorem 6 in Appendix Appendix B). The covariance structure of well-balanced 2D fBm is provided in section III.
Definition 2 (Well-balanced 2D fBm).
@@
for $s\in\mathbb{R},t\geq 0,j=1,2$, $\sigma_{1},\sigma_{2}>0$, and the constants $a^{*}_{H_{j}}$ are given by
@@
The proposed constructions of 2D fBm are closely related to the general operator fractional Brownian motion (OFBM), a family of processes even more general than vector fractional Brownian motion. OFBM is a multivariate Gaussian process defined through a matrix-valued self-similarity exponent and integral kernels acting on multidimensional Brownian noise[52, 53, 54, 55]. This connection provides helpful tools for the presented models and situates them within the broader theory of multivariate self-similar processes.
Both of our considered models coincide with OFBM with the exponent matrix $D~=~\mathrm{diag}\left(H_{1}-\frac{1}{2},\,H_{2}-\frac{1}{2}\right)$ and a kernel matrix determined by correlation parameter of the underlying noise $\rho$ and the model itself. Since in the causal 2D fBm the negative-time kernel is absent, the resulting process is inherently asymmetric in time. In contrast, the well-balanced 2D fBm corresponds to an OFBM with both positive and negative kernels present in a symmetric manner, resulting in a time-reversible process. This structural distinction is discussed in section III and is described through the asymmetry parameter $\eta_{12}$.
While OFBM theory typically imposes nontrivial admissibility conditions on the exponent and kernel matrices, our presented construction avoids these difficulties by employing correlated Brownian motions directly in the integral representation. As a consequence, the model is valid for all $H_{1},H_{2}\in(0,1)$ and $|\rho|\leq 1$ without additional constraints. Thus, the introduced 2D fBm may be viewed as a subclass of OFBM, one that naturally captures both symmetric and asymmetric cross-dependence through the choice of causal or well-balanced form of integral kernels.
@@ III.1 Covariance structure of the process
The proof of this theorem is presented in Appendix Appendix A and Appendix Appendix C. ∎
@@ IV Spectral Content
The PSD matrix plays a central role in the analysis of multidimensional stochastic processes[57]. This quantity generalizes the concept of a PSD to vector-valued processes. It is particularly valuable in multiple fields, including econometrics [58], structural engineering [59, 60], neuroscience [61], and signal processing [62], where complex systems exhibit interactions across multiple channels or variables. Since PSD is more frequently employed in the context of stationary processes, we begin this section by analyzing the increments of 2D fBm.
@@ IV.1 Spectral content of the increments
which is known as Wiener-Khinchin theorem[63]. Let us note, that for a $d$-dimensional process, the function $S_{\Delta\mathbf{Z}}$ is a $d\times d$ matrix with the diagonal elements corresponding to marginals and off-diagonal elements describing the, so called, cross power spectral density [57]. To streamline the notation, we will use now $S^{\Delta}$ whenever PSD relates to the increments process, instead of using $S_{\Delta\mathbf{Z}}$.
@@
for $j,k=1,2$. Constants $a_{H_{j}}$ and $a^{*}_{H_{j}}$ are given in Eqs. (7)) and (11), respectively. Then, for the increments of 2D fBm the power spectral density matrix $S^{\Delta}(f)$ is given by
@@
where $(x)_{+}\equiv\max\{x,0\},(x)_{-}\equiv\max\{-x,0\},D=\mathrm{diag}(H_{1},H_{2})-0.5I_{2}$, $I_{2}$ is a $2\times 2$ identity matrix, and $\overline{C}$ denotes the element-wise complex conjugate of matrix $C$. Matrix $\tilde{C}$ is equal to $C_{\textrm{c}}$ or $C_{\textrm{wb}}$ depending if we consider causal or well-balanced 2D fBm. The components $\langle S^{\Delta}_{jk}\rangle$ can be thus expressed as
@@
The proof of this theorem is presented in Appendix Appendix D. ∎
@@
The direct derivation of matrix $C$ given in Theorem 3 be found in the proof of Theorem 1 given in Appendix Appendix A and Appendix Appendix C.
@@ IV.2 Spectral content of the process
The proof of this theorem is included in Appendix Appendix E. ∎
@@
$\left.-a_{jk}\left[\frac{\cot\left(\frac{\pi}{2}(H_{j}+H_{k})\right)}{\tilde{\omega}^{H_{j}+H_{k}+1}}+\frac{2+H_{j}+H_{k}-\cos\tilde{\omega}+\sin\tilde{\omega}\cot\left(\frac{\pi}{2}(H_{j}+H_{k})\right)}{\tilde{\omega}^{H_{j}+H_{k}+2}}\right]\right\},$ (48)
@@
The proof of this theorem is based on the asymptotic expansion of the incomplete gamma function and is presented in details in Appendix Appendix F. ∎
@@
$\Re\langle S_{jk}(f,T)\rangle$ $\sim\sigma_{j}\sigma_{k}\rho_{jk}\left[\frac{1}{f^{2}}T^{H_{j}+H_{k}-1}+\frac{a_{jk}}{f^{H_{j}+H_{k}+1}}+o(1)\right],$ (49)
$\Im\langle S_{jk}(f,T)\rangle$ $\sim\sigma_{j}\sigma_{k}\eta_{jk}\left[\frac{2}{f^{2}}T^{H_{j}+H_{k}-1}-\frac{a_{jk}\cot\left(\frac{\pi}{2}(H_{j}+H_{k})\right)}{f^{H_{j}+H_{k}+1}}+o(1)\right],$ (50)
@@
$\Re\langle S_{jk}(f,T)\rangle$ $\sim\sigma_{j}\sigma_{k}\rho_{jk}\left[\frac{a_{jk}}{f^{H_{j}+H_{k}+1}}+\frac{1}{f^{2}T^{1-H_{j}-H_{k}}}+o\left(T^{H_{j}+H_{k}-1}\right)\right],$ (51)
$\Im\langle S_{jk}(f,T)\rangle$ $\sim\sigma_{j}\sigma_{k}\eta_{jk}\left[\frac{-a_{jk}\cot\left(\frac{\pi}{2}(H_{j}+H_{k})\right)}{f^{H_{j}+H_{k}+1}}+\frac{2}{f^{2}T^{1-H_{j}-H_{k}}}+o\left(T^{H_{j}+H_{k}-1}\right)\right]$ (52)
@@ V.2 Autocovariance
Figure 4: Cross-covariance function depending on $H_{1},H_{2}$ and $\rho_{12}=0.5$ for $H_{1}=H_{2}$. (a) $H_{1}=H_{2}=0.2$, (b) $H_{1}=H_{2}=0.5$, and (c) $H_{1}=H_{2}=0.7$. The solid lines represent the causal version of the model and dashed lines the well-balanced case. Markers correspond to the estimated values of the cross-covariance function; circles correspond to the causal case, squares to the well-balanced case. Here, casual and well-balanced cases overlap, as for $H_{1}=H_{2}$ the model is time-reversible regardless of the definition (cf. Eq. (15)). Increments are taken over a unit time interval, i.e., $\Delta\equiv\Delta^{\delta=1}$.
@@ V.3 Spectral content
Figure 6: Ensemble-averaged (cross-)PSD. Solid lines correspond to PSD estimated from the simulated trajectories, dashed lines to their asymptotics. (a) $\Re\langle S_{\mathbf{X},12}(f,T)\rangle$, $H_{1}=H_{2}$; (b) $\Re\langle S_{\mathbf{X},12}(f,T)\rangle$, $H_{1}\neq H_{2}$; and (c) $\Im\langle S_{\mathbf{X},12}(f,T)\rangle$, $H_{1}\neq H_{2}$. In each case, $\rho_{12}=0.5$. Since the real part for both causal and well-balanced case is the same, we plot only only one corresponding line to each set of parameteres. Conversely, the imaginary part for the well-balanced case is 0, so we plot only $\Im\langle S_{\mathbf{X},12}\rangle$ for the causal 2D fBm. In all the panels, $T=2^{16}=65,536$ and the lines are based on $N=5,000$ trajectories.
@@
In all panels, we see that the asymptotic behavior of the ensemble-averaged (cross-)PSD for increments behaves in a similar way in all cases, regardless if $H_{1}+H_{2}\gtrless 1$. The asymptotic behavior, as $f\to 0$ is proportional to $f^{1-H_{1}-H_{2}}$. Here, we see a quantitative difference of diverging or converging (cross-)PSD around 0 – if $H_{1}+H_{2}>1$ then the (cross-)PSD diverges (in 1D, it corresponds to one of the definitions of long memory [52]), while for $H_{1}+H_{2}<1$ the (cross-)PSD converges to 0.
Figure 7: Ensemble-averaged (cross-)PSD for increments. Continuous lines correspond to PSD estimated from the simulated trajectories’ increments, dashed lines to their asymptotics. (a) $\Re\langle S^{\Delta}_{12}(f,T)\rangle$, $H_{1}=H_{2}$; (b) $\Re\langle S^{\Delta}_{12}(f)\rangle$, $H_{1}\neq H_{2}$; and (c) $\Im\langle S^{\Delta}_{12}(f)\rangle$, $H_{1}\neq H_{2}$. In each case, $\rho_{12}=0.5$. Since the real part for both causal and well-balanced case is the same, we plot only only one corresponding line to each set of parameters. Conversely, the imaginary part for the well-balanced case is 0, so we plot only $\Im\langle S^{\Delta}_{12}(f)\rangle$ for the causal 2D fBm’s increments. In all the panels, $T=2^{16}=65,536$ and the lines are based on $N=5,000$ trajectories.